A Sushi is a sushi take-away chain store which offers individually packed sushi, sushi boxes and various donburi that are freshly made on-site. Last year, the company spent $135,000 hiring a marketing consultant to evaluate whether or not a new line of sushi should be launched. The consultant found that the new product would be able to generate $970,000 of additional sales revenue per year for the company. Production of the new product will involve the following activities: – A new machine has to be purchased prior to the commencement of production. The new machine will cost $2,140,000 and has a useful life of four years. For tax purposes, assume that the new machine would be fully depreciated by the straight-line method over a period of four years.

– To purchase the new machine, it appears that the company has to borrow $1,500,000 at 6% interest from its bank, resulting in additional interest expenses of $90,000 per year.

– If the company accepts this project, its annual cash expenses will increase from $5,840,000 to $6,020,000.

– The project will necessitate an increase in accounts receivable of $202,000. In addition, the marginal tax rate, the average tax rate and the required rate of return for the company are 22%, 14% and 16% respectively.

a. Determine the relevant annual after-tax cash flows associated with this project.

b. Would you accept the project if the required payback period is four years?

c. Critically discuss four problems associated with using the payback period to evaluate this
Answer #1
Amount of $135,000 spent on Marketing Consultant is sunk cost
This Cost is not relevant
a) RELEVANT ANNUAL AFTER TAX CASH FLOWS
a. Annual additional Sales Revenue $970,000
b. Increase in annual cash expense ($180,000) (5840000-6020000)
c. Annual Interest expense ($90,000)
d.=a.+b.+c. Increase in Before tax Income (excluding depreciation) $700,000
e.=d.*22% Increase in tax expenses -$154,000
f=d+e Increase in after tax Income (excluding depreciation tax shield) $546,000
g Annual Depreciation (2140000/4) $535,000
h=g*22% Annual Depreciation Tax Shield $117,700
j=f+h Annual After Tax Cash Flow $663,700
b) Payback Period :
Payback Period =Initial Cash Flow /Annual After tax Cash Flow
Initial Cash flows:
Cost of Machine ($2,140,000)
Increase in accounts receivable ($202,000)
I Net Initial Cost ($2,342,000)
Payback Period =2342000/663700                       3.53
YES, PROJECT IS ACCEPTED IF REQUIRED PAYBACK PERIOD IS 4 YEARS
c) Problems Associated with using Payback period:
1. It does not consider time value of money
2. It does not consider cash flows after the payback period
3.It does not consider wealth created by the project
4.It Rejects long gestation but profitable projects